OpenStudy (anonymous):

A square pyramid is shown. What is the surface area? A square based pyramid, with bases labeled 2.5 centimeters and side length of triangle labeled 5 centimeters. 15.25 cm2 15.625 cm2 31.25 cm2 62.625 cm2

OpenStudy (adamaero):

What is the formula for the surface area of a pyramid?

OpenStudy (danjs):

The area of the base is simply 2.5 * 2.5 TO find the area of the 4 triangles, find the height first.

OpenStudy (anonymous):

i know it now c

OpenStudy (danjs):



OpenStudy (danjs):

Find that height, then Area of triangle is Half the base times the height (0.5)*2.5*h

OpenStudy (anonymous):

Use a net to find the surface area of the right triangular prism shown below: Three rectangles next to each other with a width of 8 feet. The first and second rectangles length are unknown, and the last is 13 feet long. There are two right triangles above and below the last rectangle. One is labeled with a base of 5 feet and a height of 12 feet. 88 square feet 232 square feet 283 square feet 300 square feet

OpenStudy (danjs):

Area of Base = 2.5^2 = 6.25 cm^2 Area of 4 Triangles = 2 * (2.5) * 4.84 = 24.2 cm^3 Total Area= [6.25 + 24.2] cm^2 = 30.45 cm^2

OpenStudy (anonymous):

Use a net to find the surface area of the right triangular prism shown below: Three rectangles next to each other with a width of 8 feet. The first and second rectangles length are unknown, and the last is 13 feet long. There are two right triangles above and below the last rectangle. One is labeled with a base of 5 feet and a height of 12 feet. 88 square feet 232 square feet 283 square feet 300 square feet